# Scalable Valuation of Human Feedback through Provably Robust Model Alignment

### v1

@@ Abstract
Despite the importance of aligning language models with human preferences, crowd-sourced human feedback is often noisy—for example, preferring less desirable responses—posing a fundamental challenge to alignment. A truly robust alignment objective should yield identical model parameters even under severe label noise, a property known as redescending. We prove that no existing alignment methods satisfy this property. To address this, we propose Hölder-DPO, the first principled alignment loss with a provable redescending property, enabling estimation of the clean data distribution from noisy feedback. The aligned model estimates the likelihood of clean data, providing a theoretically grounded metric for dataset valuation that identifies the location and fraction of mislabels. This metric is gradient-free, enabling scalable and automated human feedback valuation without costly manual verification or clean validation dataset. Hölder-DPO achieves state-of-the-art robust alignment performance while accurately detecting mislabels in controlled datasets. Finally, we apply Hölder-DPO to widely used alignment datasets, revealing substantial noise levels and demonstrating that removing these mislabels significantly improves alignment performance across methods.
[…]
@@ 1 Introduction
Aligning large language models (LLMs; [1, 93, 90]) with human preferences is essential for value alignment and mitigating safety risks [63, 9]. Direct Preference Optimization (DPO; [82]) has emerged as a key method, fine-tuning LLMs using pairwise comparisons—e.g., between preferred and rejected responses [24, 17]. This framework supports alignment with a broad range of rankable values, such as helpfulness [8], toxicity reduction [50], and summarization quality [88, 95]. Human feedback is typically gathered via crowdsourcing, including from platforms like ChatGPT [1], but is often noisy—e.g., preferring undesirable responses—which can significantly deteriorate model performance [103, 23, 28]. In fact, the Anthropic HH-RLHF dataset [33] reportedly contain over 25% inconsistent feedback [86, 19, 103]. Plus, recent work [106, 41, 15] shows that models trained on smaller, high-quality datasets can outperform those trained on larger, noisier ones. However, maintaining data quality often requires costly manual review (e.g., inter-annotator agreement), highlighting the need for scalable, automated methods to assess and clean human feedback.
@@
To mitigate the impact of noisy human feedback, researchers have proposed robustifed DPO variants. Among various heuristics [73, 6, 98], two methods claim provable guarantees: Provably Robust DPO (R-DPO) [23] and Distributionally Robust DPO (Dr. DPO) [103]. However, their notion of robustness is limited to bounding the parameter estimation error—bounds that degrade as the contamination ratio $\epsilon$ (the fraction of mislabeled data) increases. In contrast, truly robust methods should estimate the clean data distribution precisely under heavy contamination. This stronger form of robustness corresponds to the redescending property [71], a classical concept in robust statistics formalized via the influence function (IF; [44, 26]). We analyze existing (robust) DPO variants through IFs and prove that none satisfy the redescending property. If this property holds, the aligned model approximates the clean data distribution—enabling principled mislabel detection, as mislabels naturally exhibit lower likelihoods being clean data points. Crucially, identifying mislabeled data has benefits beyond robustness: it supports annotator evaluation, incentivizes higher-quality feedback, and enables scalable dataset cleaning, reducing reliance on costly manual verification and promoting more transparent, interpretable training pipelines.
@@
To address this challenge, we propose Hölder-DPO, the first DPO variant with a provable redescending property. As illustrated in Figure 1, fine-tuning an LLM with Hölder-DPO not only yields a robustly aligned model but also enables the model itself to serve as a mislabel valuation metric. Unlike existing data valuation methods that rely on direct IF computation [58, 21], Hölder-DPO requires only likelihood evaluation—bypassing costly gradient-based approximations and enabling scalable dataset valuation. Hölder-DPO demonstrates superior robustness and accurately detects mislabeled examples across datasets. Applied to the Anthropic HH dataset [33], it reveals substantial noise, and removing the identified mislabels significantly improves alignment performance across methods.
@@ 2.1 Language Model Alignment by Direct Preferential Optimization
Human feedback and Bradley-Terry model. Let $x\in\mathcal{X}$ be a prompt sampled from a distribution $p(x)$ over a finite context space $\mathcal{X}$. The policy model $\pi(y\mid x)$ generates a pair of responses $(y_{1},y_{2})$. A human annotator then provides a preference between them, identifying the preferred response as $y_{\textrm{win}}$, and the rejected response as $y_{\textrm{lose}}$, i.e., $y_{\textrm{win}}\succ y_{\textrm{lose}}$. Although the underlying reward function $r^{*}(x,y)$, which ranks responses, is unobservable, Bradley–Terry (BT) model [12] provides a standard framework for modeling pairwise comparisons based solely on observed preferences:
@@
where $\sigma(a)\coloneqq 1/(1+\exp(-a))$ is the logistic function. For notational convenience, we define a pairwise data point as $s\coloneqq\{x,y_{\textrm{win}},y_{\textrm{lose}}\}$, and denote the preference dataset as $\mathcal{D}\coloneqq\{s^{(i)}\}_{i=1}^{N}$. Each data point $s^{(i)}$ is sampled independently from $p_{\mathcal{D}}(s)$, which is the following joint distribution $p_{\mathcal{D}}(s)\propto p^{*}(y_{\text{win}}\succ y_{\text{lose}}\mid x,y_{1},y_{2})\cdot\pi(y_{1},y_{2}\mid x)\cdot p(x)$.
@@
where $\pi_{\mathrm{ref}}$ is the reference model, and $\beta>0$ is a hyperparameter controlling alignment strength. The reference model is typically obtained by supervised fine-tuning (SFT) on the prompts $x$ with the preferred response $y_{\text{win}}$ only [82, 84].
The corresponding gradient is given by:
@@ 2.2 Data Contamination Model for Label Flips and Redescending Property
Let $20\leq\epsilon<1/$ 2 be contamination ratio. The generative distribution of $\widetilde{\mathcal{D}}$ under the $\epsilon$-contamination model is defined as $p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})=(1-\epsilon)p_{\mathcal{D}}(s)+\epsilon p(s_{\textrm{flip}})$, and denote $\tilde{s}$ as a contaminated data point, which can be either clean $s$ or flipped $s_{\textrm{flip}}$.
Definition 1 is widely adopted in robust statistics [47, 43, 48] and in the machine learning community [37, 32, 72]. We assume access only to the contaminated dataset $p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})$, without access to a clean validation set $p_{\mathcal{D}}(s)$—a prerequisite in many existing methods [23, 56]. As such, the contamination problem cannot be reduced to a standard classification or distribution shift task. Instead, we leverage a tail contamination assumption to enable estimation of the contamination ratio $\epsilon$:
@@
Intuitively, if we had access to the ground-truth clean-data likelihood $\sigma(g_{\theta^{*}}(s))$, it would assign high values to clean data points and negligibly small values to flipped ones. This principle is well-established in robust statistics [31, 52], and similar assumptions are implicitly adapted in prior work [55, 13, 4, 108]. Moreover, this assumption unlocks heavy contamination scenarios, in contrast to typical assumptions that consider only infinitesimally small perturbations, i.e., $\epsilon\approx 0$ [40, 59, 21, 62, 61, 107, 51]. This is particularly important for preference datasets, which may contain over 20% noisy labels.
Redescending property. We evaluate robustness to the data contamination through the learned LLM parameters. Let $\theta^{*}$ denote the optimal parameters learned from the clean dataset $p_{\mathcal{D}}$, and $\theta^{*}(\epsilon)$ denote those learned from the $\epsilon$-contaminated dataset $p^{(\epsilon)}_{\widetilde{\mathcal{D}}}$. Intuitively, the impact of contamination can be measured by $\|\theta^{*}(\epsilon)-\theta^{*}\|$, and its first-order approximation, known as IF [43], is commonly used:
$\mathrm{IF}(x,\theta,p_{\mathcal{D}})\coloneqq\bigg{|}\frac{\partial\theta^{*}(\epsilon)}{\partial\epsilon}\bigg{|}_{\epsilon=0}\bigg{|}.$ (4)
@@
$\lim_{r_{\theta}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\mathrm{IF}(x,\theta,p_{\mathcal{D}})=0.$
In the above, the limit $r_{\theta}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty$ corresponds to the worst-case scenario for DPO, where the model is forced to learn the least preferable $y_{\text{lose}}$ as best due to label flip, i.e., $y_{\mathrm{lose}}^{\textrm{flip}}=y_{\mathrm{win}}$. Intuitively, if the IF remains zero even when this adversarial setting, the learning objective can be considered robust, as contamination does not affect the learned parameters, i.e., $\theta^{*}(\epsilon)\approx\theta^{*}$ [^1] Up to higher-order approximation error $\mathcal{O}(\epsilon^{2})$. Definition 2 can be seen as an extension of the redescending property [71]—a classical concept in robust statistics.
@@ 2.3 Related work
Robust DPOs and RLHFs. A broad range of DPO variants have been proposed to improve the robustness of the alignment objective [73, 6, 98, 23, 103]. Reinforcement-learning-based approaches also constitute a popular family of alignment methods [25, 8, 76, 54], with several recent works introducing robustified objectives [69, 14, 80]. However, none provides theoretical guarantees for the redescending property or for detecting mislabeled data. We prove their limitations in Section 3.
Dataset valuation and cleaning. Dataset valuation has been studied using IFs [40, 59, 21, 62, 61, 107, 51], Shapley values [100], and approximate unrolling techniques [7]. Various heuristic methods have also been proposed for identifying and filtering mislabeled data, including prompt-based filtering [74, 89, 65, 101, 67], attribution-based approaches [99, 35, 11, 77], and perplexity-based filtering [56]. In contrast, our method is the first to provide theoretical guarantees for dataset filtering under heavy data contamination.
@@ 3 Existing DPO Variants Are Neither Robust Nor Able to Identify Mislabels
Let $\theta^{*}$ be the optimum learnt from the clean dataset $p_{\mathcal{D}}$, and $\theta^{*}(\epsilon)$ learnt from the $\epsilon$-contaminated dataset $p^{(\epsilon)}_{\widetilde{\mathcal{D}}}$. Let $\mathcal{L}_{\mathrm{gen}}(\pi_{\theta};\pi_{\mathrm{ref}})$ be a generic DPO loss function corresponding to a specific DPO variant. Assume that the Hessian $\nabla_{\theta}^{2}\mathcal{L}_{\textrm{gen}}(s,\pi_{\theta})|_{\theta=\theta^{*}}$ is positive definite. Then, the $s_{\textrm{flip}}$-dependent component of the IF for this DPO variants is given by:
$\mathrm{IF}(x,\theta,p_{\mathcal{D}})\propto|\mathbb{E}_{p(s_{\textrm{flip}})}[\nabla_{\theta}\mathcal{L}_{\mathrm{gen}}(s_{\textrm{flip}},\pi_{\theta^{*}})]|,$ (5)
@@
DPO [82] $\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(g_{\theta}(\tilde{s}))]$
IPO [6] $\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}\bigg{[}\bigg{(}\sigma\bigg{(}\frac{g_{\theta}(\tilde{s})}{\beta}\bigg{)}-\frac{1}{2\beta}\bigg{)}^{2}\bigg{]}$
C-DPO [73] $(1-c)\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(g_{\theta}(\tilde{s}))]-c\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(-g_{\theta}(\tilde{s}))]$
R-DPO [23] $\frac{1-c}{1-2c}\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(g_{\theta}(\tilde{s}))]-\frac{c}{1-2c}\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(-g_{\theta}(\tilde{s}))]$
Dr. DPO [103] $-\beta^{\prime}\log\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}\bigg{[}\exp\bigg{(}\frac{\log\sigma(g_{\theta}(\tilde{s}))}{\beta^{\prime}}\bigg{)}\bigg{]}$
@@
See Appendices E and G for full proofs and detailed expressions of the IFs for each DPO variant. Figure 2 illustrates Theorem 3, showing how the weight term in each IF, which governs robustness, behaves under extreme perturbations. Both DPO and Dr. DPO exhibit the same asymptotic trend, with $\lim_{r_{\theta}(x,y^{\text{flip}}_{\text{lose}})\rightarrow\infty}\text{IF}=a$, where $a$ is a non-zero constant. R-DPO shifts this limit upward with a constant $c$, while C-DPO pushes it downward, but neither approaches zero. In contrast, our proposed Hölder-DPO, introduced in the next section, completely nullifies the influence of outliers, demonstrating strong robustness (redescending), not merely boundedness.

### v2

@@ Abstract
Despite the importance of aligning language models with human preferences, crowd-sourced human feedback is often noisy—for example, preferring less desirable responses—posing a fundamental challenge to alignment. A truly robust alignment objective should yield identical model parameters even under severe label noise, a property known as redescending. We prove that no existing alignment methods satisfy this property. To address this, we propose Hölder-DPO, the first principled alignment loss with a provable redescending property, enabling estimation of the clean data distribution from noisy feedback. The aligned model estimates the likelihood of clean data, providing a theoretically grounded metric for dataset valuation that identifies the location and fraction of mislabels. This metric is gradient-free, enabling scalable and automated human feedback valuation without costly manual verification or clean validation dataset. Hölder-DPO achieves state-of-the-art robust alignment performance while accurately detecting mislabels in controlled datasets. Finally, applied to Anthropic HH-RLHF dataset, it reveals substantial noise levels and removing these mislabels significantly improves alignment performance across methods. The code is available [^1] https://github.com/ma921/HolderDPO.
@@ 1 Introduction
Aligning large language models (LLMs; [1, 95, 92]) with human preferences is essential for value alignment and mitigating safety risks [65, 9]. Direct Preference Optimization (DPO; [84]) has emerged as a key method, fine-tuning LLMs using pairwise comparisons—e.g., between preferred and rejected responses [25, 18]. This framework supports alignment with a broad range of rankable values, such as helpfulness [8, 51], and summarization quality [90, 97]. Human feedback is typically gathered via crowdsourcing (e.g., ChatGPT [1]), but is often noisy—e.g., preferring undesirable responses—which can significantly deteriorate model performance [105, 24, 29]. In fact, the Anthropic HH-RLHF dataset [34] reportedly contain over 25% inconsistent feedback [88, 20, 105]. Plus, recent work [107, 42, 15] shows that models trained on smaller, high-quality datasets can outperform those trained on larger, noisier ones. However, maintaining data quality often requires costly manual review (e.g., inter-annotator agreement), highlighting the need for scalable, automated assessment methods.
@@
To mitigate the impact of noisy human feedback, researchers have proposed robustifed DPO variants. Among various heuristics [76, 6, 100], two methods claim provable guarantees: Provably Robust DPO (R-DPO) [24] and Distributionally Robust DPO (Dr. DPO) [105]. However, their notion of robustness is limited to bounding the parameter estimation error—bounds that degrade as the fraction of mislabelled data $\epsilon$ increases. In contrast, truly robust methods should estimate the clean data distribution precisely under heavy contamination. This stronger form of robustness corresponds to the redescending property [74]—the influence of extreme outliers diminishes to zero—a concept in robust statistics defined via the influence function (IF; [45, 27]). We analyze existing DPO variants through IFs and prove that none satisfy the redescending property. If this property holds, the aligned model approximates the clean data distribution—enabling principled mislabel detection, as mislabels naturally exhibit lower likelihoods being clean data points. Crucially, identifying mislabelled data has benefits beyond robustness: it supports annotator evaluation, incentivises higher-quality feedback, and enables scalable dataset cleaning, reducing reliance on costly manual verification.
@@
To address this challenge, we propose Hölder-DPO, the first alignment loss with a provable redescending property, stably learning clean data distribution even under heavy contamination (e.g., $\epsilon=0.4$). Furthermore, our loss also enables the aligned model itself to serve as a mislabel valuation metric (see Figure 1). Ours is a first-of-its-kind principled metric that provides a theoretically grounded estimator of both the likelihood and fraction of (mis-)labelled data. Furthermore, our method is highly scalable—it avoids costly gradient-based computations required by prior work [60, 22] and eliminates the need for manual verification [58], thereby enabling fully automated feedback cleaning. Applied to the Anthropic HH dataset [34], it reveals substantial noise ($\epsilon\approx 0.25$), and removing the identified mislabels significantly improves alignment performance across methods.
@@ 2.1 Language Model Alignment by Direct Preferential Optimization
Human feedback and Bradley-Terry model. Let $x\in\mathcal{X}$ be a prompt sampled from a distribution $p(x)$ over a finite context space $\mathcal{X}$. The policy model $\pi(y\mid x)$ generates a pair of responses $(y_{1},y_{2})$. A human annotator then provides a preference between them, identifying the preferred response as $y_{\textrm{win}}$, and the rejected response as $y_{\textrm{lose}}$, i.e., $y_{\textrm{win}}\succ y_{\textrm{lose}}$. Although the underlying reward function $r^{*}(x,y)$, which ranks responses, is unobservable, Bradley–Terry (BT) model [12] provides a standard framework for modelling pairwise comparisons based solely on observed preferences:
@@
where $\sigma(a)\coloneqq 1/(1+\exp(-a))$ is the logistic function. We define a pairwise data point as $s\coloneqq\{x,y_{\textrm{win}},y_{\textrm{lose}}\}$, and denote the preference dataset as $\mathcal{D}\coloneqq\{s^{(i)}\}_{i=1}^{N}$. Each data point $s^{(i)}$ is sampled independently from $p_{\mathcal{D}}(s)$, where $p_{\mathcal{D}}(s)\propto p^{*}(y_{\text{win}}\succ y_{\text{lose}}\mid x,y_{1},y_{2})\cdot\pi(y_{1},y_{2}\mid x)\cdot p(x)$.
@@
where $\pi_{\mathrm{ref}}$ is the reference model, and $\beta>0$ is a hyperparameter controlling alignment strength. The reference model is typically obtained by supervised fine-tuning (SFT) on the prompts $x$ with the preferred response $y_{\text{win}}$ only [84, 86]. The corresponding gradient is given by:
@@ 2.2 Data Contamination Model for Label Flips and Redescending Property
Let $0\leq\epsilon<1/2$ be contamination ratio. The generative distribution of $\widetilde{\mathcal{D}}$ under the $\epsilon$-contamination model is defined as $p^{(\epsilon)}_{\widetilde{\mathcal{D}}}(\tilde{s})=(1-\epsilon)p_{\mathcal{D}}(s)+\epsilon p(s_{\textrm{flip}})$, and denote $\tilde{s}$ as a contaminated data point, which can be either clean $s$ or flipped $s_{\textrm{flip}}$.
[…]
@@
Intuitively, if we had access to the ground-truth clean-data likelihood $\sigma(g_{\theta^{*}}(s))$, it would assign high values to clean data points and negligibly small values to flipped ones. This principle is well-established in robust statistics [32, 53], and implicitly adapted in prior work [57, 13, 4, 109]. Moreover, this assumption unlocks heavy contamination scenarios, in contrast to typical assumptions that consider only infinitesimally small perturbations, i.e., $\epsilon\approx 0$ [41, 61, 22]. This distinction is particularly important for noisy human preference datasets (e.g., $\epsilon\approx 0.25$ for HH-RLHF [88, 20]).
Redescending property. We evaluate robustness to the data contamination through the learned LLM parameters. Let $\theta^{*}$ denote the optimal parameters learned from the clean dataset $p_{\mathcal{D}}$, and $\theta^{*}(\epsilon)$ denote those learned from the $\epsilon$-contaminated dataset $p^{(\epsilon)}_{\widetilde{\mathcal{D}}}$. Intuitively, the impact of contamination can be measured by $\theta^{*}(\epsilon)-\theta^{*}$, and its first-order approximation, known as IF [44], is commonly used:
$\mathrm{IF}(s_{\textrm{flip}},\theta,p_{\mathcal{D}})\coloneqq\frac{\partial\theta^{*}(\epsilon)}{\partial\epsilon}\bigg|_{\epsilon=0}.$ (4)
@@
$\lim_{\hat{r}_{\theta}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty}\|\mathrm{IF}(s_{\textrm{flip}},\theta,p_{\mathcal{D}})\|=0,$
where $\|\cdot\|$ denotes the Euclidean norm.
The limit $\hat{r}_{\theta}(x,y_{\mathrm{lose}}^{\textrm{flip}})\to\infty$ occurs when the data point $s_{\textrm{flip}}$ attains the lowest reward, i.e., $g_{\theta}(s_{\textrm{flip}})\rightarrow$-$\infty$. Yet, the label flip forces the model to learn this least preferable sample as best, representing the most adversarial data-poisoning scenario that a single label flip can induce. Intuitively, if the IF remains zero even under this adversarial condition, the learning objective can be considered robust: the effect of the worst possible data point on the model parameters vanishes, i.e., $\theta^{*}(\epsilon)\approx\theta^{*}$ [^2] Up to higher-order approximation error $\mathcal{O}(\epsilon^{2})$.
@@ 2.3 Related work
Robust DPOs and RLHFs. A broad range of DPO variants have been proposed to improve the robustness of the alignment objective [76, 6, 100, 24, 105]. Reinforcement-learning-based approaches also constitute a popular family of alignment methods [26, 8, 78, 55], with several recent works introducing robustified objectives [72, 14, 82]. However, none provides theoretical guarantees for the redescending property or for detecting mislabelled data. We prove their limitations in Section 3.
Dataset valuation and cleaning. Dataset valuation has been studied using IFs [41, 61, 22, 64, 63, 108, 52], Shapley values [102], and unrolling [7]. Various heuristic methods have also been proposed for identifying and filtering mislabelled data, including prompt-based filtering [77, 91, 68, 103, 70], attribution-based approaches [101, 36, 11, 79], and perplexity-based filtering [58]. In contrast, our method is the first to provide theoretical guarantees for dataset filtering under heavy contamination.
@@ 3 Existing DPO Variants Are Neither Robust Nor Able to Identify Mislabels
Let $\theta^{*}$ be the optimum learnt from the clean dataset $p_{\mathcal{D}}$, and $\theta^{*}(\epsilon)$ learnt from the $\epsilon$-contaminated dataset $p^{(\epsilon)}_{\widetilde{\mathcal{D}}}$. Let $\mathcal{L}_{\mathrm{gen}}(\pi_{\theta};\pi_{\mathrm{ref}})$ be a generic DPO loss function corresponding to a specific DPO variant. Assume that the Hessian $\nabla_{\theta}^{2}\mathcal{L}_{\textrm{gen}}(s,\pi_{\theta})|_{\theta=\theta^{*}}$ is positive definite. [^3] This is a local assumption around the optimal parameters $\theta^{*}$. Such local assumptions are standard in IF analysis for non-convex deep learning models, as global convexity rarely holds (e.g., [56, 22]). Then, the $s_{\textrm{flip}}$-dependent component of the IF for this DPO variants is given by:
$\mathrm{IF}(s_{\textrm{flip}},\theta,p_{\mathcal{D}})\propto\mathbb{E}_{p(s_{\textrm{flip}})}[\nabla_{\theta}\mathcal{L}_{\mathrm{gen}}(s_{\textrm{flip}},\pi_{\theta^{*}})],$ (5)
@@
DPO [84] $)]\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(g_{\theta}(\tilde{s}$))]
IPO [6] $\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}\bigg[\bigg(\frac{g_{\theta}(\tilde{s})}{\beta}-\frac{1}{2\beta}\bigg)^{2}\bigg]$
C-DPO [76] $)](1-c)\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(g_{\theta}(\tilde{s}))]-c\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(-g_{\theta}(\tilde{s}$))]
R-DPO [24] $)]\frac{1-c}{1-2c}\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(g_{\theta}(\tilde{s}))]-\frac{c}{1-2c}\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}[-\log\sigma(-g_{\theta}(\tilde{s}$))]
Dr. DPO [105] $\beta^{\prime}\log\mathbb{E}_{p^{(\epsilon)}_{\widetilde{\mathcal{D}}}}\bigg[\exp\bigg(\frac{\log\sigma(g_{\theta}(\tilde{s}))}{\beta^{\prime}}\bigg)\bigg]$
@@
See Appendices E and G for full proofs and detailed expressions of the IFs for each DPO variant. We now provide an intuitive interpretation of Theorem 3. The IF can be viewed as the gradient of the generic objective $\mathcal{L}_{\text{gen}}$ (see Eq. (5)), which can be decomposed into two components. For example, in the vanilla DPO case (Eq. (3)), we have:
$\nabla_{\theta}\mathcal{L}(s,\pi_{\theta})=\underbrace{\sigma(-g_{\theta}(s))}_{\textrm{IF weights.}}\,\,\underbrace{g_{\text{grad}}(s)}_{\textrm{direction}}.$
Here, the gradient term $g_{\text{grad}}(s)$ determines the direction of the update, while the IF weight $\sigma(-g_{\theta}(s))$ acts as a scaling factor that controls the strength of the update for each data point. If the IF weight tends to zero in the worst-case scenario (i.e., $g_{\theta}(s_{\textrm{flip}})\rightarrow$-$\infty$), the method is robust, as it effectively ignores highly confident but misaligned (flipped) samples. Figure 2 visualises the behavior of IF weights across different objectives under extreme reward perturbations. Both DPO and Dr. DPO exhibit the same asymptotic trend, with $\lim_{r_{\theta}(x,y^{\text{flip}}_{\text{lose}})\rightarrow\infty}\text{IF}=a$, where $a$ is a non-zero constant. R-DPO shifts this limit upward with a constant $c$, while C-DPO pushes it downward, but neither approaches zero. In contrast, our proposed Hölder-DPO, introduced in the next section, completely nullifies the influence of outliers, demonstrating strong robustness (redescending), not merely boundedness.